# MoodAngels: A Retrieval-augmented Multi-agent Framework for Psychiatry Diagnosis

### v1

@@ F.2.4 Privacy Evaluation
F.2.5 Logistic Detection Evaluation
@@
F.3 Usage Declaration
This work and its contents, including the synthetic mental health diagnostic dataset and diagnostic agent, are provided for academic and research purposes only. None of the material constitutes medical, clinical, or therapeutic advice. The synthetic data is artificially generated and does not represent real patient information. No warranties, express or implied, are offered regarding the accuracy, reliability, or suitability of the dataset, agent, or any related content. The authors and contributors are not responsible for any errors, omissions, or consequences arising from the use of this repository.
F.4 IRB approvals
Figure 7 displays the IRB approval document alongside its English translation. To comply with double-blind review requirements, all identifying information (e.g., hospital and applicant details) has been redacted.
Figure 7: Institutional Review Board (IRB) Approval: Original and English Translation.
Appendix G Agent Details
G.1 Symptom Matching
G.2 Prompts of Agents
G.2.1 Prompts of Angel.R, Angel.D, and Angel.C
Role:
You are a psychiatry diagnosis expert.
Objective:
Diagnose whether the visitor {digit_id} has a mood disorder (including depression and bipolar disorder), which is characterized by depressive or manic symptoms.
Constraints:
1.You may only use the following actions.
2.You must act proactively. Always keep this in mind when planning actions.
Available Actions:
Three varients of angels are different in actions. When using a specific variant, please comment out the other two.
Angel.R
toggle_visitor_record: Retrieves structured medical records for the visitor, if available. Also extracts top-5 symptoms related to DSM-5 diagnostic criteria. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"}
get_scale_performances: Retrieves the visitor’s performance on the top 5% most mood disorder-related questions from psychological scales. The correlation score represents the statistical relevance to mood disorders. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"}
Finish: Completes the diagnosis process. Args:{"answer": ‘yes’ or ‘no’ (indicating whether the visitor has a mood disorder)", "reasons": "A summary of key reasons supporting the decision."}
Angel.D
toggle_visitor_record: Retrieves structured medical records for the visitor, if available. Also extracts top-5 symptoms related to DSM-5 diagnostic criteria. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"}
get_scale_performances: Retrieves the visitor’s performance on the top 5% most mood disorder-related questions from psychological scales. The correlation score represents the statistical relevance to mood disorders. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"}
previous_cases_display: If the visitor has medical record information, this tool compares the case with past structured medical records in the database and extracts the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: "name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"
previous_scales_display: This tool compares the visitor’s scale performance with past case scale performances in the database, and extracts the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: "name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"
Finish: Completes the diagnosis process. Args:"answer": ‘yes’ or ‘no’ (indicating whether the visitor has a mood disorder)", "reasons": "A summary of key reasons supporting the decision."
Angel.C
toggle_visitor_record: Retrieves structured medical records for the visitor, if available. Also extracts top-5 symptoms related to DSM-5 diagnostic criteria. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"}
get_scale_performances: Retrieves the visitor’s performance on the top 5% most mood disorder-related questions from psychological scales. The correlation score represents the statistical relevance to mood disorders. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"}
previous_cases_analysis: If the visitor has medical record information, this tool compares the case with past structured medical records in the database, extracts and analyzes the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: "name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"
previous_scales_analysis: This tool returns a comparison of the visitor’s scale performance with past case scale performances in the database, extracts and analyzes the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: "name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"
Finish: Completes the diagnosis process. Args:"answer": ‘yes’ or ‘no’ (indicating whether the visitor has a mood disorder)", "reasons": "A summary of key reasons supporting the decision."
Resources:
You are a large language model trained on vast textual data, including factual medical knowledge.
Best Practices:
1. Always refer to similar cases (medical records or scale data) for diagnosis. However, do not over-rely on past diagnoses due to the personalized nature of mental health conditions.
2. Consider all available information (medical records & scale performance). Always examine similar records (if available) and similar scale data for reference. Do not rely solely on one data source.
3 .Prioritize clinical evaluation over self-reported scales in case of contradictions. Self-reported terms like "occasionally", "sometimes", and "frequently" are subjective, making it difficult for visitors to assess their condition accurately.
4. Only consider symptoms relevant to mood disorders. The visitor may have other mental disorders that are not classified as mood disorders.
5. Mood disorders include only depression and bipolar disorder, characterized by depressive or manic symptoms.
Response Format:
Generate a JSON string following the structure below. Do not include any extra text or explanation. {"action": {"name": "action name", "args": {"args name": "args value"}}, "thoughts": {"plan": "Briefly describe the list of short-term and long-term plans", "criticism": "Constructive self-criticism", "observation": "Summary of the current step returned to the user", "reasoning": "Reasoning behind the decision"}}
G.2.2 Prompts of Two Debate Agents
Role:
Two sides of the debate. When using a specific one, please comment out the other.
the positive side
You are a psychiatry diagnosis expert. You believe the current visitor has a mood disorder. Present your argument and persuade the opposing side.
the negative side
You are a psychiatry diagnosis expert. You believe the current visitor does not have a mood disorder. Present your argument and persuade the opposing side.
Visitor Information:
This is a challenging case, and three agents have already provided their diagnoses:
Agent without prior case retrieval believes the visitor {is/isn’t, diagnosis by Angel.R} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.R}
Agent who retrieved past cases but did not analyze them believes the visitor {is/isn’t, diagnosis by Angel.D} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.D}
Agent who retrieved and analyzed past cases believes the visitor {is/isn’t, diagnosis by Angel.C} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.C}
Original Case Information:
Visitor’s Medical Record: {The visitor’s medical record in Step 3}
Visitor’s Scale Performances: {The visitor’s performances on selected scales}
Reference from Past Medical Records: {Top-5 similar medical records with analysis}
Reference from Past Scale Performances: {Top-5 similar scale performances with analysis}
Constraints:
Stick to your stance. Every case presented for discussion is inherently controversial, meaning your viewpoint is strongly justified.
You cannot interact with physical objects. If an action is absolutely necessary, request the user to perform it. If the user refuses and no alternative exists, terminate the process to avoid wasting time and effort.
Resources:
You are a large language model trained on extensive textual data, including factual medical knowledge.
All relevant visitor information is provided. Make full use of it.
Best Practices:
Consider all available visitor information (medical records & scale performances). Examine both similar medical records and similar scale performances as references. Do not base conclusions on a single source.
Due to the personalized nature of psychiatric diagnosis, do not overly rely on similar case diagnoses.
Self-reported questionnaire results may contradict clinician evaluations. In such cases, prioritize the clinician’s assessment. Terms like "occasionally," "sometimes," and "frequently" in self-reports are subjective and may lead to misinterpretation of the visitor’s condition.
Only symptoms related to mood disorders should influence your diagnosis. The visitor may have other mental disorders that are not classified as mood disorders.
Mood disorders include only depression and bipolar disorder, characterized by depressive or manic symptoms.
Response Format:
Generate a JSON string in the following format. Do not include any extra text or explanation.
{"response": "Your reasoning for why the visitor has (or does not have) a mood disorder, along with a rebuttal to the opposing argument.", "thoughts": {"plan": "Briefly outline short-term and long-term plans", "criticism": "Constructive self-criticism"}}
G.2.3 Prompts of the Judge
Role:
You are a psychiatry diagnosis expert acting as the judge in this consultation. Your role is to determine whether each debate round should conclude and make the final diagnosis.
Visitor Information:
This is a challenging case, and three agents have already provided their diagnoses:
Agent without prior case retrieval believes the visitor {is/isn’t, diagnosis by Angel.R} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.R}
Agent who retrieved past cases but did not analyze them believes the visitor {is/isn’t, diagnosis by Angel.D} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.D}
Agent who retrieved and analyzed past cases believes the visitor {is/isn’t, diagnosis by Angel.C} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.C}
Original Case Information:
Visitor’s Medical Record: {The visitor’s medical record in Step 3}
Visitor’s Scale Performances: {The visitor’s performances on selected scales}
Reference from Past Medical Records: {Top-5 similar medical records with analysis}
Reference from Past Scale Performances: {Top-5 similar scale performances with analysis}
Constraints:
Fulfill your role as a judge and base your decision solely on the arguments presented by both sides.
You cannot interact with physical objects. If an action is absolutely necessary to complete the task, request the user to perform it. If the user refuses and no alternative exists, terminate the process to avoid wasting time and effort.
Resources:
You are a large language model trained on extensive textual data, including factual medical knowledge.
All relevant visitor information is provided. Make full use of it.
Best Practices:
Consider all available visitor information (medical records & scale performances). Examine both similar medical records and similar scale performances as references. Do not base conclusions on a single source.
Due to the personalized nature of psychiatric diagnosis, do not overly rely on similar case diagnoses.
Self-reported questionnaire results may contradict clinician evaluations. In such cases, prioritize the clinician’s assessment. Terms like "occasionally," "sometimes," and "frequently" in self-reports are subjective and may lead to misinterpretation of the visitor’s condition.
Only symptoms related to mood disorders should influence your diagnosis. The visitor may have other mental disorders that are not classified as mood disorders.
Mood disorders include only depression and bipolar disorder, characterized by depressive or manic symptoms.
Response Format:
Generate a JSON string in the following format. Do not include any extra text or explanation.
{"judge": "Do you believe the debate should end? Answer only ’yes’ or ’no’.", "diagnose": "Do you believe the visitor has a mood disorder? Answer only ’yes’ or ’no’.", "thoughts": { "plan": "Briefly outline short-term and long-term plans.", "criticism": "Constructive self-criticism.", "judge_reasons": "Your reasoning for why the debate should or should not end.", "reasoning": "Your reasoning for whether the visitor does or does not have a mood disorder." }}
Appendix H Prompt Settings in Experiments
For prompts in bare LLMs, we employ the key element combination format derived from medical records (Step 3 in Figure 3) along with the overall performance metrics (total score and corresponding descriptions) of each available scale. It is important to note that some scales do not rely on a total score to assess performance; instead, they use individual questions or specific question combinations to evaluate a particular dimension (e.g., CTQ scale and MCCB tests). Additionally, certain scales include not only multiple-choice questions but also open-ended questions, making it impossible to calculate a total score (e.g., NSSI scale). However, since several scales closely related to mood disorders utilize a fully multiple-choice format with a total score calculation, the information provided to the baselines remains comprehensive. A prompt template with an example is shown in Figure 8.
@@ Appendix H Prompt Settings in Experiments
Appendix I Case Study

### v2

@@ D.2.4 Privacy Evaluation
D.2.5 Logistic Detection Evaluation
@@
Appendix E Experiment
E.1 Experimental Settings
Baseline Prompts. For prompts in bare LLMs, we employ the key element combination format derived from medical records (Step 3 in Figure 3) along with the overall performance metrics (total score and corresponding descriptions) of each available scale. It is important to note that some scales do not rely on a total score to assess performance; instead, they use individual questions or specific question combinations to evaluate a particular dimension (e.g., CTQ scale and MCCB tests). Additionally, certain scales include not only multiple-choice questions but also open-ended questions, making it impossible to calculate a total score (e.g., NSSI scale). However, since several scales closely related to mood disorders utilize a fully multiple-choice format with a total score calculation, the information provided to the baselines remains comprehensive. A prompt template with an example is shown in Figure 8.
@@ E.1 Experimental Settings
Retrieval number. Table 18 evaluates how retrieval number $k$ affects model performance, and $k=5$ yields the optimal balance across evaluation metrics. Note that the optimal value of $k$ may vary across datasets and should be adjusted accordingly when using the MoodAngels diagnostic framework.
Table 18: Performances with different retrieval numbers.
Setting ACC MCC
Angel.Dk=5 0.923 0.837
Angel.Dk=3 0.906 0.794
Angel.Dk=1 0.923 0.733
Angel.R(k=0) 0.920 0.829
E.2 Experimental Results
E.2.1 Human Evaluation of Generated Reasons
The debate agents generate reasons for their position by points. We evaluated their performances by three licensed psychiatry experts on our co-author clinician’s team. Metrics are clinical coherence (0-2 points), information completeness (0-2 points), and clarity of expression (0-2 points), as detailed in Table 19. The assessment results are shown in Table 20.
Table 19: Evaluation criteria of generated reasons.
Clinical Coherence: whether reasoning aligns with clinical logic and evidence.
2 = Fully coherent and professionally sound;
1 = Minor logical flaws;
0 = Major errors or conflicts.
Information Completeness: inclusion of key diagnostic elements (e.g., symptoms, scales, course, exclusions, impairment).
2 = Comprehensive and sufficient;
1 = Missing important elements;
0 = Lacks key diagnostic info.
Clarity and Expression: clarity and professionalism of the explanation.
2 = Clear and professionally worded;
1 = Some ambiguity or terminology misuse;
0 = Unclear or confusing.
Table 20: Evaluation results of generated reasons.
Model coh1 com1 clarity1 coh2 com2 clarity2 coh3 com3 clarity3 avg_coh avg_com avg_clarity
GPT-4o 0.5 0.9 0.7 0.2 0.5 0.9 0.6 1 1.1 0.43 0.5 0.9
Angel.R 0.9 1.1 0.9 0.6 0.7 1.2 1 1.3 1.3 0.87 1.03 1.17
Angel.D 0.8 1.2 0.9 0.6 0.6 1 0.9 1.3 1.2 0.83 1.03 0.73
Angel.C 0.9 1.2 1 0.7 0.8 1.1 1.1 1.4 1.3 1.27 1.4 1.13
multi-Angels 1.1 1.3 1.1 0.8 0.8 1.1 1.1 1.1 1.1 1.3 1.3 1.1
We report Fleiss’ Kappa as the measure of inter-annotator agreement (IAA), and the results are shown in Table 21. The results demonstrate the high quality of our generated outputs. The higher inter-annotator agreement suggests that the outputs of our multi-agent system are more consistent, clinically aligned, and better recognized by experts.
Table 21: Evaluation IAA of generated reasons.
Model coherence completeness clarity
GPT-4o 0.6 0.26 0.13
Angel.R 0.55 0.37 0.19
Angel.D 0.69 0.28 0.25
Angel.C 0.69 0.33 0.27
multi-Angels 0.67 0.45 0.54
E.2.2 Subgroup Performance Analysis
We evaluate the performance of Angel.R across demographic subgroups on the patient subset, containing 190 cases of mood disorder and 56 cases of other diseases. The performances are shown in Table 22.
For gender subgroups, despite a nominal accuracy difference ($\Delta=0.088$), the p-value (0.083 > 0.05) and a comparable MCC difference ($\Delta=0.018$) suggest no statistically significant gender bias. The slightly higher female accuracy likely reflects dataset characteristics rather than model bias.
For age subgroups, no significant accuracy difference (p=0.110) is observed between adults and children, with MCC values differing by only 0.048. It indicates that our model generalizes well across age groups without age-specific calibration.
Table 22: Subgroup Performances.
Group ACC MCC
All Patients 0.845 0.489
Gender Subgroups
Male 0.784 0.466
Female 0.872 0.484
Age Subgroups
Adult 0.795 0.500
Female 0.873 0.452
E.3 Ablation Study
In this ablation study, we evaluate the robustness of our framework by changing input order and temperature.
Input Order Sensitivity. We use the random.shuffle() function to randomize the input order during the debate stage, and the results are shown in Table 23.
Table 23: Performances with different input orders.
Setting ACC MCC
multi-Angels (ordered input) 0.925 0.834
multi-Angels (shuffled input) 0.927 0.838
Temperature Sensitivity. We vary the temperature of GPT-4o to assess the output stability of Angel.R, and the results are shown in Table 24.
Table 24: Performances with different LLM temperatures.
Setting ACC MCC
Angel.Rtemp=1(default) 0.920 0.829
Angel.Rtemp=0.7 0.922 0.836
Angel.Rtemp=0.3 0.923 0.836
Angel.Rtemp=0 0.911 0.808
These results indicate that variations in input order and temperature have minimal impact on diagnostic performance, demonstrating the stability and robustness of our framework.
E.4 Case Study